# Retrieval-Enhanced Knowledge Editing for Multi-Hop Question Answering in Language Models

### v1

@@ 6. Conclusion
We propose a novel LLM editing framework for multi-hop QA. We employ a strategy that utilizes mutual information maximization for the retrieval of facts and a self-optimizing technique for pruning redundant information. This approach effectively addresses the challenges of integrating real-time knowledge updates in LLMs. Theoretical justifications and comprehensive evaluations demonstrate RAE’s effectiveness in enhancing the accuracy of updated LLM responses. Our contributions mark a substantial advancement in model editing research, presenting a promising direction for the integration of dynamic knowledge in language models.
@@ Appendix A Theoretical Justification
Consider latent concepts $\theta_{c}$ is inferred from a knowledge graph $G_{S}$, and a question $q$ is then generated based on this concept. (For instance, this can be as straightforward as prompting LLMs to formulate a question about a given set of facts). The following Markov chain: $G_{S}\rightarrow\theta_{c}\rightarrow q$ exists, indicating that $q$ is conditionally independent to $G_{s}$. According to the chain rule for mutual information, we can have:
@@ Appendix B Implementation Details
In our fact retrieval process, we employ a beam search with a width of 2, meaning we consider only two candidate relations at each hop. For the total number of retrieval hops $n$, we set it to 4, 5, and 6 for 2-hop, 3-hop, and 4-hop questions, respectively. For simplicity, we omit the term $p(q|h_{1})$ as it remains constant for each question, and its exclusion does not impact empirical performance. Regarding the SEARC baseline, we pair smaller models with larger auxiliary models: GPT-2-small (124M) (Radford et al., 2019) with GPT-2-xl (1.3B) (Radford et al., 2019), GPT-2-large (774M) (Radford et al., 2019) with GPT-J (6B) (Wang and Komatsuzaki, 2021), and GPT-Neo (2.7B) (Black et al., 2021) with both Falcon (7B) (Almazrouei et al., 2023) and Vicuna (7B) (Chiang et al., 2023), as well as Llama2 (7B) (Touvron et al., 2023a). We implement these models with the code and checkpoints available from Huggingface library (Wolf et al., 2019).
Acknowledgements.

### v2

@@ 6. Conclusion
We propose a novel LLM editing framework for multi-hop QA that employs mutual information maximization for fact retrieval and a self-optimizing technique to prune redundant data. This effectively tackles the integration of real-time knowledge updates in LLMs. Our extensive evaluations confirm the framework’s ability to enhance the accuracy of LLM responses, marking significant progress in model editing and dynamic knowledge integration research.
Acknowledgements.
@@ Appendix A Theoretical Justification
Consider latent concepts $\theta_{c}$ is inferred from a knowledge graph $G_{S}$, and a question $q$ is then generated based on this concept. The following Markov chain: $G_{S}\rightarrow\theta_{c}\rightarrow q$ exists, indicating that $q$ is conditionally independent to $G_{s}$. According to the chain rule for mutual information, we can have:
@@ Appendix B Implementation Details
In our fact retrieval process, we employ a beam search with a width of 2, meaning we consider only two candidate relations at each hop. For the total number of retrieval hops $n$, we set it to 4, 5, and 6 for 2-hop, 3-hop, and 4-hop questions, respectively. For simplicity, we omit the term $p(q|h_{1})$ as it remains constant for each question, and its exclusion does not impact empirical performance. Regarding the SEARC baseline, we pair smaller models with larger auxiliary models: GPT-2-small (124M) (Radford et al., 2019) with GPT-2-xl (1.5B) (Radford et al., 2019), GPT-2-large (774M) (Radford et al., 2019) with GPT-J (6B) (Wang and Komatsuzaki, 2021), and GPT-Neo (2.7B) (Black et al., 2021) with both Falcon (7B) (Almazrouei et al., 2023) and Vicuna (7B) (Chiang et al., 2023), as well as Llama2 (7B) (Touvron et al., 2023a). We implement these models with the code and checkpoints available from Huggingface library (Wolf et al., 2019). For Subgraph Retriever (Zhang et al., 2022), we first retrieve the knowledge graph, and then we incorporate it into the editing LLMs with an editing prompt.